OpenStudy (anonymous):

how do i solve this system: 5x+2y =7 -x/2 + y/6 = 3/2

OpenStudy (amistre64):

solve for y in the top: y = -5x/2 +7/2 plug it into the bottom equation: -x/2 + (-5x/2 +7/2)/6 = 3/2 Solve for x: -x/2 -5x/12 +7/12 = 3/2 -6x/12 -5x/12 +7/12 = 18/12 -6x -5x +7 =18 -11x = 11 x=-1 Now that we have a solid number for x; plug it into the equation we found for y: y = -5(-1)/2 +7/2 y= 5/2 + 7/2 y= 12/2 = 6 Double check for accuracy :)

OpenStudy (anonymous):

how do you know that the 7 only goes to where the y is.. i thoguth you were supposed to plug 0 in for one of the values and then solve and then use that answer for the equatiuon